# Turning Tabular Foundation Models into Graph Foundation Models

### v1

@@ Abstract
While foundation models have revolutionized such fields as natural language processing and computer vision, their application and potential within graph machine learning remain largely unexplored. One of the key challenges in designing graph foundation models (GFMs) is handling diverse node features that can vary across different graph datasets. Although many works on GFMs have been focused exclusively on text-attributed graphs, the problem of handling arbitrary features of other types in GFMs has not been fully addressed. However, this problem is not unique to the graph domain, as it also arises in the field of machine learning for tabular data. In this work, motivated by the recent success of tabular foundation models like TabPFNv2, we propose G2T-FM, a simple graph foundation model that employs TabPFNv2 as a backbone. Specifically, G2T-FM augments the original node features with neighborhood feature aggregation, adds structural embeddings, and then applies TabPFNv2 to the constructed node representations. Even in a fully in-context regime, our model achieves strong results, significantly outperforming publicly available GFMs and performing on par with well-tuned GNNs trained from scratch. Moreover, after finetuning, G2T-FM surpasses well-tuned GNN baselines, highlighting the potential of the proposed approach. More broadly, our paper reveals a previously overlooked direction of utilizing tabular foundation models for graph machine learning tasks. [^1] Our source code is available at https://github.com/yandex-research/G2T-FM.
@@ 1 Introduction
Despite their remarkable success in such areas as computer vision and natural language processing, the development of foundation models for graph data has been less advanced. The challenges of developing graph foundation models (GFMs) stem from the fact that graphs are not actually a single domain, but rather a way to represent data from different domains. These domains use graphs to represent very different structures, such as social networks, web networks, road networks, co-purchasing networks, molecules, connectomes, or even abstract objects and their relations. Thus, successful GFMs should be able to work with graphs from different domains representing very different objects with nodes and very different relations with edges, which is a rather formidable task that requires overcoming many serious challenges. Two key challenges faced by GFMs are the ability to transfer to new feature spaces and target spaces. Graphs from different domains often have different node features and different targets, making it difficult to design GFMs that can work across various types of graphs. Some existing GFMs restrict themselves to text-attributed graphs (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024), which allows them to use pretrained text encoders. Another approach is to use simple dimensionality reduction methods like SVD and PCA (Xia & Huang, 2024; Zhao et al., 2024a; Wang et al., 2025a; Yu et al., 2025), which allow transforming all feature spaces to a space with a fixed predefined number of features. But these approaches do not allow for fully and effectively leveraging arbitrary node features in graphs from new domains.
@@
However, we note that the challenges of transferring to new feature and target spaces are not exclusive to graphs. Tabular data — one of the most widespread data modalities in machine learning — is similar to graph-structured data in that it does not constitute a single domain but is rather a way to represent data from different domains, and different tabular datasets come with different feature and target spaces. Thus, tabular foundation models face similar issues to GFMs. While tabular foundation models are not as developed as foundation models for language or vision, they have seen increased interest recently (Van Breugel & Van Der Schaar, 2024), and the first successful approaches have been proposed (Hollmann et al., 2023; 2025; Mueller et al., 2025; Ma et al., 2024; Qu et al., 2025). For instance, TabPFNv2 (Hollmann et al., 2025) demonstrates strong performance in both in-context learning and finetuning scenarios, and it has recently gained significant attention from the community.
@@
We argue that GFM developers should take inspiration from tabular foundation model approaches, as they have to deal with many of the same problems. In this paper, we take a first step in this direction and show that tabular foundation models, such as TabPFNv2 (Hollmann et al., 2025), can be effectively adapted to graph datasets. We introduce a simple framework named Graph-to-Table Foundation Model (G2T-FM), which transforms graph tasks into tabular ones and solves them with a tabular foundation model. More specifically, we augment the original features with neighborhood feature aggregations (Bazhenov et al., 2025), classical structure-based features (node degree, PageRank, and the eigenvectors of the graph Laplacian), and learnable structure-based encodings (Kanatsoulis et al., 2025). Then, we apply TabPFNv2 to the constructed node representations.
@@
Our empirical results indicate that this straightforward framework achieves strong results in a fully in-context regime, significantly outperforming existing publicly available GFMs and performing on par with well-tuned GNNs trained from scratch. Moreover, after finetuning, G2T-FM surpasses well-tuned GNN baselines, highlighting the potential of the proposed approach and the positive transfer brought by the usage of foundation models.
@@
  • We draw attention to a promising and previously overlooked direction of applying tabular foundation models to graph machine learning.
  • As a proof of concept, we introduce G2T-FM, a simple framework that uses a tabular foundation model (TabPFNv2 in our experiments) as the backbone of a graph foundation model.
@@ 2.1 Graph Foundation Models for Node Classification
In this work, we primarily focus on node-level tasks such as node classification and node regression. While many GFMs are limited to text-attributed graphs (TAGs) (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024), they are not truly general as graphs in many domains come with other feature types that are difficult to describe with text. Because of that, we specifically discuss methods that can be applied to graphs with arbitrary numerical and categorical node features which frequently appear in various real-world industrial applications.
@@
Some graph foundation models use self-supervised learning (SSL) objectives to guide their pretraining process (Zhao et al., 2024a; Yu et al., 2025; Wang et al., 2025a), whereas others employ supervised learning strategies (Lachi et al., 2024; Finkelshtein et al., 2025). Notably, several works (Xia et al., 2024; Xia & Huang, 2024) reduce the node classification task to link prediction. Specifically, for each label in a downstream task, they create a virtual node that is connected to all train nodes of that class. After that, node classification reduces to prediction of links to those virtual class nodes, and these models are pretrained on link prediction task.
@@
One of the key challenges for graph foundation models is handling heterogeneous features that can vary significantly across different datasets. Some approaches address this by focusing exclusively on text-attributed graphs (TAGs), sometimes additionally converting all features of other types to text, and then applying a text encoder (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024). Methods aiming to deal with arbitrary features often rely on simple dimensionality reduction techniques such as SVD or PCA to obtain feature embeddings (Xia & Huang, 2024; Zhao et al., 2024a; Wang et al., 2025a; Yu et al., 2025). There are alternatives, such as learning dimension encoding modules that produce feature transformations (Zhao et al., 2024b) or replacing node attribute values with their statistical dependencies (Shen et al., 2025), but they seem to be less popular. We also highlight Finkelshtein et al. (2025), which constructs separate embeddings for each (node, feature) pair, enabling a more fine-grained representation of feature information.
@@
Handling the structure is relatively easier, as graph neural networks are particularly well-suited for this task, and they are inherently capable of processing arbitrary graph structures. Consequently, many GFMs simply adopt GNNs as their backbone to handle graph structure (Zhao et al., 2024a; Finkelshtein et al., 2025; Yu et al., 2025; Wang et al., 2025a). In addition to GNN-based approaches, some methods use matrix decomposition techniques such as SVD applied to graph-derived matrices (for example, the normalized adjacency matrix or the sum of its powers), to encode structural information (Xia et al., 2024). However, while GNNs can in principle operate on any graph, their performance may still be limited due to varying graph structures. To address this, some works implement additional mechanisms specifically designed to handle structural differences (Yu et al., 2025; Wang et al., 2025a).
@@ 2.2 Limitations of Existing GFMs
Many existing graph foundation models are specifically designed for text-attributed graphs, where nodes or edges have associated textual information (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024). These models typically leverage large language models or other text encoders to process textual attributes, integrating natural language representations with graph structures. While this approach can be effective for certain domains such as academic networks or knowledge graphs, it limits the application of GFMs to a broader range of graphs where such text attributes are not available. For instance, for graphs representing transportation networks, biological networks, or transaction networks (commonly used for fraud detection tasks), which often come with rich numerical and categorical features, the reliance exclusively on textual information restricts the model’s usability and effectiveness. As a result, many current GFMs may not generalize well to graphs with non-textual attributes, hindering their adoption across diverse real-world scenarios.
@@
Most publicly available GFMs are designed and evaluated on classification tasks, where the goal is to predict categorical labels for nodes, edges, or entire graphs. To date, none of the popular GFMs besides TS-GNN (Finkelshtein et al., 2025) support regression tasks, where the output is a continuous value rather than a class label. This is a substantial limitation because many important graph-based applications require regression instead of classification. The lack of support for regression tasks reduces the practical applicability of current GFMs and highlights an important area for future research.
@@
Some recent studies on graph foundation models have described their methods as operating in a “zero-shot” setting (Xia & Huang, 2024; Xia et al., 2024). Typically, these approaches introduce virtual nodes that represent target classes and connect them to the corresponding real nodes with known class labels. Then, the node classification problem reduces to predicting links between the test nodes and the appropriate virtual nodes. This process makes it possible to perform evaluation on unseen graphs without additional finetuning. While inventive and interesting, this technique does not truly realize zero-shot learning. Strictly speaking, zero-shot learning means that no labeled examples of the target classes are available during evaluation. However, the described method requires labeled nodes to be connected to virtual class nodes for effective link prediction. Therefore, the correct term for this setup should be “in-context learning”, since evaluation does not involve further finetuning but still depends on the access to labeled training samples.
This inconsistency in terminology may lead to misleading comparisons with baseline approaches. For instance, the aforementioned studies (Xia & Huang, 2024; Xia et al., 2024) compare their “zero-shot” performance against the one-shot and five-shot results of other baselines, yet they do not clearly report the number of training samples used in “zero-shot” evaluation of the proposed method, which makes the comparison harder to interpret.
@@ 2.3 Tabular Foundation Models
The field of tabular foundation models (TFMs) was pioneered by the TabPFN model (Hollmann et al., 2023) that was designed to address any tabular problem off-the-shelf. TabPFN employs a transformer-like architecture and works in the in-context learning regime, with the entire downstream training set serving as the prompt. The pretraining of TabPFN was performed on a large number of synthetic datasets designed to mimic typical tabular tasks. The more recent model, TabPFNv2 (Hollmann et al., 2025), employs a more powerful backbone architecture, pretraining on a broader spectrum of synthetic datasets, and advanced techniques of data preprocessing. Nowadays, new TFMs are emerging regularly (Mueller et al., 2025; Ma et al., 2024; Qu et al., 2025; Zhang & Robinson, 2025), and their success is exploited beyond the domain of pure tabular tasks, e.g, for time series forecasting (Hoo et al., 2025). In our work, we demonstrate that TFMs can also serve as a core building block for graph foundation models.
@@ 3 Graph-to-Table Foundation Model
However, these challenges are not unique to graph machine learning, and they also arise in tabular machine learning. Recent advances in tabular deep learning, such as the development of foundation models like TabPFNv2 (Hollmann et al., 2025), offer promising solutions to the problems of working with diverse feature and target spaces.
Figure 1: Overview of the proposed G2T-FM.
We argue that tabular foundation models can be used to create better graph foundation models, and, in particular, they can help handling different feature and target spaces. As a proof of concept, we introduce Graph-to-Table Foundation Model (G2T-FM), which addresses the challenge of learning on graphs with diverse and heterogeneous node features and different targets. Figure 1 provides an overview of our method.
To process heterogeneous node features, we employ the TabPFNv2 model (Hollmann et al., 2025). This model is designed for tabular data only, so to make it applicable to graph-structured data, we introduce a graph-based preprocessing step that encodes the information about graph into the node features. Our goal is to capture the information about different aspects of the graph structure as well as the interplay between the graph structure and the node features. Hence, the new augmented feature vector consists of the original node features and the following graph-based components.
@@
We also include basic node structural characteristics — node degree, PageRank score and Laplacian eigenvectors. First two are classic node centrality measures that indicate how “important” a particular node is, however, they do it in very different ways. The node degree captures strictly local information, while PageRank also captures global information. Then, we compute the first $K$ eigenvectors of the graph Laplacian and consider the corresponding $K$-dimensional embeddings as additional feature vectors. Laplacian embeddings encode the information about the position of a node in graph relative to other nodes. Thus, such node representations provide valuable information that supplements the centrality measures.
@@
These encodings have been proposed in Kanatsoulis et al. (2025). The basic idea of PEARL is to generate a random value for each node and then apply a GNN using these values as node features. Such random initialization increases the expressive power of GNNs by breaking the structural symmetries. This procedure is repeated $M$ times (each with new random node features) and the resulting node embeddings are averaged so that node permutation equivariance still holds in the limit. Kanatsoulis et al. (2025) propose using this as a learnable module, so that the encodings are trained to improve downstream performance. However, we noticed that PEARL encodings are also useful without training, i.e., when we use a randomly initialized GNN to obtain the node embeddings. Thus, this module can produce both non-learnable and learnable representations, where non-learnable representations do not require optimizing (training) any parameters.
These new features allow us to provide the tabular foundation model with information about many properties of the graph underlying the given dataset. We concatenate them with the original node attributes and use the resulting augmented feature vector as an input to TabPFNv2. The resulting model can be applied in a fully in-context regime and, as we show below, it performs on par with GNNs that are well-tuned for a single dataset. In finetuning regime, we jointly tune PEARL and TabPFNv2.
As a final remark, we discuss the symmetries of G2T-FM. Following Finkelshtein et al. (2025), a graph foundation model should satisfy three natural inductive biases: (i) feature permutation invariance; (ii) label permutation equivariance; and (iii) node permutation equivariance. G2T-FM satisfies these symmetries in distribution: its outputs depend on internal randomness, but the distribution of the outputs remains invariant or equivariant under the corresponding permutations. See Appendix B for details.

### v2

@@ Abstract
While foundation models have revolutionized such fields as natural language processing and computer vision, their potential in graph machine learning remains largely unexplored. One of the key challenges in designing graph foundation models (GFMs) is handling diverse node features that can vary across different graph datasets. While many works on GFMs have focused exclusively on text-attributed graphs, the problem of handling arbitrary features of other types in GFMs has not been fully addressed. However, this problem is not unique to the graph domain, as it also arises in the field of machine learning for tabular data. In this work, motivated by the recent success of tabular foundation models (TFMs) like TabPFNv2 or LimiX, we propose G2T-FM, a simple framework for turning tabular foundation models into graph foundation models. Specifically, G2T-FM augments the original node features with neighborhood feature aggregation, adds structural embeddings, and then applies a TFM to the constructed node representations. Even in a fully in-context regime, our model achieves strong results, significantly outperforming publicly available GFMs and performing competitively with, and often better than, well-tuned GNNs trained from scratch. Moreover, after finetuning, G2T-FM surpasses well-tuned GNN baselines. In particular, when combined with LimiX, G2T-FM often outperforms the best GNN by a significant margin. In summary, our paper reveals the potential of a previously overlooked direction of utilizing tabular foundation models for graph machine learning tasks. [^1] Our source code is available at https://github.com/yandex-research/G2T-FM.
@@ 1 Introduction
Despite their remarkable success in such areas as computer vision and natural language processing, the development of foundation models for graph data has been less advanced. The challenges of developing graph foundation models (GFMs) stem from the fact that graphs are not actually a single domain, but rather a way to represent data from different domains. These domains use graphs to represent very different structures, such as social networks, web networks, road networks, co-purchasing networks, molecules, connectomes, or even abstract objects and their relations. Thus, successful GFMs should be able to work with graphs from different domains representing very different objects with nodes and very different relations with edges, which is a rather formidable task that requires overcoming many serious challenges. Two key challenges faced by GFMs are the ability to transfer to new feature spaces and target spaces. Graphs from different domains often have different node features and different targets, making it difficult to design GFMs that can work across various types of graphs. Some existing GFMs restrict themselves to text-attributed graphs (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024), which allows them to use pretrained text encoders. Another approach is to use simple dimensionality reduction methods like SVD and PCA (Xia & Huang, 2024; Zhao et al., 2024a; Wang et al., 2025a; Yu et al., 2025), which allow transforming all feature spaces to a space with a fixed predefined number of features. However, these approaches do not allow for fully and effectively leveraging arbitrary node features in graphs from new domains.
@@
The challenges of transferring to new feature and target spaces are not, however, exclusive to graphs. Tabular data — one of the most widespread data modalities in machine learning — is similar to graph-structured data in that it does not constitute a single domain but is a way to represent data from different domains. Thus, tabular datasets come with different feature and target spaces, so tabular foundation models face similar issues to GFMs. While tabular foundation models are not as developed as foundation models for language or vision, they have seen increased interest recently (Van Breugel & Van Der Schaar, 2024), with the first successful approaches being proposed (Hollmann et al., 2023; 2025; Mueller et al., 2025; Ma et al., 2024; Qu et al., 2025). For instance, TabPFNv2 (Hollmann et al., 2025) demonstrates strong performance in both in-context and finetuning regimes, and it has recently gained significant attention from the community.
@@
This parallel suggests that developers of GFMs can draw inspiration from tabular foundation models, as they have to deal with many of the same challenges. In this paper, we take a first step in this direction and show that tabular foundation models, such as TabPFNv2 (Hollmann et al., 2025) and LimiX (Zhang et al., 2025), can be effectively adapted to graph datasets. We introduce a simple framework named Graph-to-Table Foundation Model (G2T-FM), which transforms graph tasks into tabular ones and solves them with a tabular foundation model. More specifically, we augment the original features with neighborhood feature aggregations (Bazhenov et al., 2025), classical structure-based features (node degree, PageRank, and the eigenvectors of the graph Laplacian), and learnable structure-based encodings (Kanatsoulis et al., 2025). Then, we apply a tabular foundation model to the constructed node representations to get predictions.
@@
Our empirical results indicate that this straightforward framework achieves strong results in a fully in-context regime, significantly outperforming existing publicly available GFMs and performing competitively with, and often better than, well-tuned GNNs trained from scratch. Moreover, after finetuning, G2T-FM surpasses well-tuned GNN baselines, with especially strong improvements obtained when G2T-FM uses LimiX as a tabular foundation model (see Table 2). These results highlight the potential of the proposed approach and the positive transfer brought by the usage of foundation models.
@@
  • We identify a promising and previously overlooked direction of applying tabular foundation models to graph machine learning.
  • As a proof of concept, we introduce G2T-FM, a simple framework that uses a tabular foundation model (TabPFNv2 and LimiX in our experiments) as the backbone of a graph foundation model.
@@ 2.1 Graph foundation models for node classification
In this work, we primarily focus on node-level tasks such as node classification and node regression. While many GFMs are limited to text-attributed graphs (TAGs) (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024), they are not truly general as graphs in many domains involve non-textual features. Therefore, we specifically discuss methods applicable to graphs with arbitrary numerical and categorical node features which frequently appear in various real-world industrial applications.
@@
Some graph foundation models use self-supervised learning (SSL) objectives to guide their pretraining process (Zhao et al., 2024a; Yu et al., 2025; Wang et al., 2025a), whereas others employ supervised learning strategies (Lachi et al., 2024; Finkelshtein et al., 2025). Notably, several works (Xia et al., 2024; Xia & Huang, 2024) reduce the node classification task to link prediction. Specifically, for each label in a downstream task, they create a virtual node that is connected to all training nodes of that class. Node classification thus becomes a task of predicting links to those virtual class nodes, for which the models are pretrained.
@@
One of the key challenges for graph foundation models is handling heterogeneous features that can vary significantly across different datasets. Some approaches address this by focusing exclusively on text-attributed graphs (TAGs), sometimes additionally converting non-textual features to text, and then applying a text encoder (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024). Methods aiming to deal with arbitrary features often rely on simple dimensionality reduction techniques such as SVD or PCA to obtain feature embeddings (Xia & Huang, 2024; Zhao et al., 2024a; Wang et al., 2025a; Yu et al., 2025). There are alternatives, such as learning dimension encoding modules that produce feature transformations (Zhao et al., 2024b) or replacing node attribute values with their statistical dependencies (Shen et al., 2025), but these appear less common. We also highlight Finkelshtein et al. (2025), which constructs separate embeddings for each (node, feature) pair, enabling a more fine-grained representation of feature information.
@@
Handling the structure is more straightforward, as graph neural networks are particularly well-suited for this task, and they are inherently capable of processing arbitrary graph structures. Consequently, many GFMs simply adopt GNNs as their backbone to handle graph structure (Zhao et al., 2024a; Finkelshtein et al., 2025; Yu et al., 2025; Wang et al., 2025a). In addition to GNN-based approaches, some methods use matrix decomposition techniques such as SVD applied to graph-derived matrices (for example, the normalized adjacency matrix or the sum of its powers), to encode structural information (Xia et al., 2024). However, while GNNs can in principle operate on any graph, their performance may still be limited due to varying graph structures. To address this, some works implement additional mechanisms specifically designed to handle structural differences (Yu et al., 2025; Wang et al., 2025a).
@@ 2.2 Limitations of existing GFMs
Many existing graph foundation models are specifically designed for text-attributed graphs, where nodes or edges have associated textual information (Wang et al., 2024b; He & Hooi, 2024; Liu et al., 2024). These models typically leverage large language models or other text encoders to process textual attributes, integrating natural language representations with graph structures. While this approach can be effective for certain domains such as academic networks or knowledge graphs, it limits the applicability of GFMs across a broader range of graphs where such text attributes are not available. For instance, for graphs representing transportation networks, biological networks, or transaction networks (commonly used for fraud detection tasks), which often come with rich numerical and categorical features, the reliance exclusively on textual information restricts the model’s usability and effectiveness. As a result, many current GFMs may not generalize well to graphs with non-textual attributes, hindering their adoption across diverse real-world scenarios.
@@
Most publicly available GFMs are designed and evaluated on classification tasks, where the goal is to predict categorical labels for nodes, edges, or entire graphs. To date, no popular GFMs, aside from TS-GNN (Finkelshtein et al., 2025), support regression tasks, where the output is a continuous value rather than a class label. This is a substantial limitation because many important graph-based applications require regression instead of classification. The lack of support for regression tasks reduces the practical applicability of current GFMs and highlights an important area for future research.
@@
Some recent studies on graph foundation models have described their methods as operating in a “zero-shot” setting (Xia & Huang, 2024; Xia et al., 2024). Typically, these approaches introduce virtual nodes that represent target classes and connect them to the corresponding real nodes with known class labels. Then, the node classification problem reduces to predicting links between the test nodes and the appropriate virtual nodes. This process makes it possible to perform evaluation on unseen graphs without additional finetuning. While inventive and interesting, this technique does not truly realize zero-shot learning. Strictly speaking, zero-shot learning means that no labeled examples of the target classes are available during evaluation. However, the described method requires labeled nodes to be connected to virtual class nodes for effective link prediction. Therefore, the correct term for this setup should be “in-context learning”, since evaluation does not involve further finetuning but still depends on access to labeled training samples. This inconsistency in terminology may lead to misleading comparisons with baseline approaches. For instance, the aforementioned studies (Xia & Huang, 2024; Xia et al., 2024) compare their “zero-shot” performance against the one-shot and five-shot results of other baselines, yet they do not clearly report the number of training samples used in “zero-shot” evaluation of the proposed method, which makes the comparison harder to interpret.
@@ 2.3 Tabular foundation models
The field of tabular foundation models (TFMs) was pioneered by the TabPFN model (Hollmann et al., 2023) that was designed to address any tabular problem off-the-shelf. TabPFN employs a transformer-like architecture and works in the in-context learning regime, with the entire downstream training set serving as the prompt. The pretraining of TabPFN was performed on a large number of synthetic datasets designed to mimic typical tabular tasks. The more recent model, TabPFNv2 (Hollmann et al., 2025), employs a more powerful backbone architecture, pretraining on a broader spectrum of synthetic datasets, and advanced techniques of data preprocessing. Nowadays, new TFMs are emerging regularly (Mueller et al., 2025; Ma et al., 2024; Qu et al., 2025; Zhang & Robinson, 2025; Zhang et al., 2025), and their success is exploited beyond the domain of pure tabular tasks, e.g., for time series forecasting (Hoo et al., 2025). In our work, we demonstrate that TFMs can also serve as a core building block for graph foundation models.
@@ 3 Graph-to-Table Foundation Model
However, these challenges are not unique to graph machine learning, and they also arise in tabular machine learning. Recent advances in tabular deep learning, such as the development of foundation models like TabPFNv2 (Hollmann et al., 2025), offer promising solutions for handling diverse feature and target spaces.
Figure 1: Overview of the proposed G2T-FM framework.
We argue that tabular foundation models can be used to create better graph foundation models, and, in particular, they can help handle different feature and target spaces. As a proof of concept, we introduce Graph-to-Table Foundation Model (G2T-FM) framework, which addresses the challenge of learning on graphs with diverse and heterogeneous node features and different targets. Figure 1 provides an overview of our method.
To process heterogeneous node features, we employ a tabular foundation model like TabPFNv2 (Hollmann et al., 2025) or LimiX (Zhang et al., 2025). TFMs are designed for tabular data only, so to make them applicable to graph-structured data, we introduce a graph-based preprocessing step that encodes graph information into the node features. Our goal is to capture the information about different aspects of the graph structure as well as the interplay between the graph structure and the node features. Hence, the new augmented feature vector consists of the original node features and the following graph-based components.
@@
We also include basic node structural characteristics — node degree, PageRank score, and Laplacian eigenvectors. The first two are classic node centrality measures that indicate how “important” a particular node is; however, they do so in different ways. The node degree captures strictly local information, while PageRank also captures global information. Then, we compute the first $K$ eigenvectors of the graph Laplacian and consider the corresponding $K$-dimensional embeddings as additional feature vectors. Laplacian embeddings encode a node’s position within the graph relative to other nodes. Thus, such node representations provide valuable information that supplements the centrality measures.
@@
These encodings have been proposed in Kanatsoulis et al. (2025). The basic idea of PEARL is to generate a random value for each node and then apply a GNN using these values as node features. Such random initialization increases the expressive power of GNNs by breaking the structural symmetries. This procedure is repeated $M$ times (each with new random node features) and the resulting node embeddings are averaged so that node permutation equivariance still holds in the limit. Kanatsoulis et al. (2025) propose using this as a learnable module, so that the encodings are trained to improve downstream performance. However, we noticed that PEARL encodings are also useful without training, i.e., when we use a randomly initialized GNN to obtain node embeddings. Thus, this module can produce both non-learnable and learnable representations, where non-learnable representations do not involve parameter optimization (training).
These new features allow us to provide the tabular foundation model with information about many properties of the graph underlying the given dataset. We concatenate them with the original node attributes and use the resulting augmented feature vector as an input to TFM. The resulting model can be applied in a fully in-context regime and, as we show below, it performs on par with GNNs that are well-tuned for a single dataset. In the finetuning regime, we jointly tune PEARL and TFM.
As a final remark, we discuss the symmetries of G2T-FM. Following Finkelshtein et al. (2025), a graph foundation model should satisfy three natural inductive biases: (i) feature permutation invariance; (ii) label permutation equivariance; and (iii) node permutation equivariance. In Appendix C, we show that G2T-TabPFNv2 satisfies these symmetries in distribution: its outputs depend on internal randomness, but the distribution of the outputs remains invariant or equivariant under the corresponding permutations.